CHAPTER 16 (3.)

In 2018, Ryan Management collected rent revenue for 2019 tenant occupancy. For financial reporting, the rent is recorded as deferred revenue and then recognized as income in the period tenants occupy rental property, but for income tax reporting it is taxed when collected. The deferred portion of the rent collected in 2018 was $65 million. Taxable income is $190 million. No temporary differences existed at the beginning of the year, and the tax rate is 40%. Suppose the deferred portion of the rent collected was $55 million at the end of 2019. Taxable income is $210 million.
  
Prepare the appropriate journal entry to record income taxes.
Answer #1
For 2018:
Debit Credit
Income tax expense 50
Deferred tax asset 26 =65*40%
       Income tax payable 76 =190*40%
For 2019:
Debit Credit
Income tax expense 88
       Deferred tax asset 4 =(65-55)*40%
       Income tax payable 84 =210*40%